# Optimizing Bayesian model selection for equation of state of cold neutron stars

### v2

@@ I Introduction
There are four principal approaches in the literature for Bayesian model selection of the dense matter EOS. In the first approach, evidence for different EOSs is obtained by directly measuring appropriately chosen physical parameters, such as their polytropic [25] or spectral [26] indices and the sound speed [27, 28, 29]. In the second approach, EOSs are compared against each other by computing the evidence for EOS models represented by - curves [30, 31, 32, 33]. For instance, Markakis et al. [34] provide correlations using numerical relativity simulations of the inspiral and postmerger GW signals from BNS mergers that can be used to constrain the EOS. Agathos et al. [30], among the first to employ a Bayesian pipeline for EOS model selection, computed pairwise Bayes factors for each EOS and a collection of simulated BNS events. However, they approximated the tidal deformability of NSs by a Taylor expansion around a fiducial mass.
@@
In the present work, we follow the model selection techniques from Ref. [31] and apply them to simulated BNS events for future GW observatories with improved sensitivities. Moreover, we calculate and compare the evidence for EOSs in different subspaces of the 4-dimensional parameter space characterized by the component masses and corresponding tidal deformabilities ---. This exhaustive comparison allows us to determine the optimal strategy for selecting the true EOS using the fourth approach.
Given data which a model should explain, Bayesian evidence penalizes complex models and models with more parameters, technically known as Occam’s penalty. There are two known ways of enforcing Occam’s penalty. In the first case, one favors a model that conforms to a subjective prior before performing Bayesian inference. The other approach is naturally included in Bayesian inference, where the computation of evidence in a higher-dimensional space, corresponding to models with greater number of parameters, leads to a lower evidence [40, 41]. We incorporate the EOS-dependent maximum mass limits by reweighting the posterior distribution, which is initially agnostic to EOS, assuming an arbitrarily large prior maximum mass [42].
@@
We defer the study of the impact of eccentricity future analyses. Roy and Saini 56 has talked about the impact of eccentricity on the tidal deformability and concluded that it could have a large impact for BNS events observed by the XG detectors when the signal is analyzed from 10 Hz. We assume NSs to have finite, but small, spins which implies spin-induced quadrupole moments are sub-dominant effects in the waveform models. We have not looked at the impact of waveform uncertainties on our model selection results. We refer to Refs. [57, 43, 58, 20] for such discussions. Reali et al. 59 presents biases due to astrophysical confusion noise and found significant biases even for high signal-to-noise ratio (SNR) BNS mergers observed in next-generation detectors (however, also see Gupta et al. 60 which finds the effect of foreground noise on parameter estimation to be minimal). It has also been argued that NS oscillations with timescales of the order of binary period could lead to biased estimate of NS tidal deformability and hence the EOS [61, 62, 63]. They mention the relative bias to be maximum of 50% for intermediate mass NS which is within the statistical uncertainty for current and near future configurations of GW detectors.
@@ II.1 Bayesian Evidence Calculation using EOS-agnostic PE results
where is the dimensionality of the prior parameter space which four our case is 4 for the case of BNS mergers. The total number of parameter for GW inference of BNS mergers has 17 parameters but, in our calculations, we start with the posterior PDF which is marginalized along every parameters except two masses () and two tidal deformabilities ().
@@
Due to our application of a narrow prior range (required by the relative binning formalism), the range of evidence integration for each event is much smaller than the range of the EOS-agnostic PE run, except for events near the maximum mass of the EOS model. Hence, the factor is different from 1.0 only for small samples of events ().
@@ II.2 Combining Evidence from multiple events
where is the EOS-specific component mass prior correction factor taken into account differently for each EOS. We take this to be a uniform distribution between and the maximum mass allowed by that specific EOS. The factor is there to undo the prior used for the calculation of evidence for each event with the following value: for the calculation of 3D and 4D evidence and for the calculation of 2D evidence in the subspace. We keep our integration range slightly outside the range used in Bayesian PE run (except for the samples beyond the maximum mass of EOS), to allow for the kernel density estimation (KDE) to keep a conservative view of EOS-specific prior correction factor.
@@ II.3.1 Choice of EOS
Formally, to compare two models, it is essential to measure the difference between them, which has traditionally been referred to by the names of EOSs, highlighting the fundamental assumptions in the construction of EOSs. Although such concepts of distinguishability have significance, they are not useful in quantifying their differences. We formally quantify the disparity between different EOSs as the distance between them (refer to Eqn. (14) of Kashyap et al. [55]), with the caveat that this metric does not entirely capture their differences. In addition, other distance measures and correlations with populations will be crucial considerations, which we aim to explore in our upcoming research.
@@ III.2 Distribution of evidence
In the space, there are different ways in which can be obtained from when computing the evidence. Therefore, degeneracies are anticipated to represent a fundamental limitation in differentiating distinct EOS models, as two different - relationships may yield identical observed posterior data. We find that the ambiguity in model selection due to the broad posterior distributions (due to the large noise in O5 and A) is greater than the apparent problem of degeneracy described above. Hence, we expect that will not be able to distinguish models in future GW detectors (such as ECC). However, we find that such differences are negligible and are important only when we combine the evidence from several events. As seen in Fig. LABEL:fig:BF-O5-ECC, the degeneracy in is shown where two EOSs compete with the injected EOS, SLy in this case, up to a large number of events. However, the subspace evidence does slightly better.
@@ III.3 Cumulative Evidence from multiple BNS events
Additional clarity on detector sensitivity can be deduced from cumulative evidence plots. For example, the Adetector configuration is more capable of distinguishing a cluster of EOS that are similar to each other but significantly different from the injected EOS compared to O5 (Fig. LABEL:fig:logevi-O5-ECC). In the case of optimistic ECC observatories, the evidence for some of the (noninjected) EOS models are numerically zero, as shown in Fig. LABEL:fig:L2vsBF-etalamt and Fig. LABEL:fig:L2vsBF-mlam2D-4D.
@@ III.4 Criterion for distinguishability and Method Efficacy
Moreover, we find that ignoring the pairwise correlation between and significantly improves the selection of the correct model. This is because such correlations do not contain information about the EOS and ignoring this i.e. avoid using the joint posterior, gives us the evidence in a smaller parameter space where Bayesian evidence calculations do not suffer from Occam’s penalty. This approach offers a new lens through which we can view and understand the interconnections between different methods proposed in the literature on model selection using the parameter .
@@ IV Conclusions and Summary
For detectors with O5 sensitivity, evidence calculations across parameter space reveal no meaningful differences between the models in question. However, for the Aand ECC networks, performing evidence calculations in the subspace - is more effective in identifying the correct model than in other spaces. This is because excluding the mass ratio from evidence calculation does not eliminate any details concerning the EOS of nuclear matter. Additionally, this choice reduces the dimensional complexity of the evidence calculation, thereby effectively avoiding Occam’s penalty.

### v3

@@ I Introduction
There are four principal approaches in the literature for Bayesian model selection of the dense matter EOS. In the first approach, evidence for different EOSs is obtained by directly measuring appropriately chosen physical parameters, such as their polytropic [25] or spectral [26] indices and the sound speed [27, 28, 29]. In the second approach, EOSs are compared against each other by computing the evidence for EOS models represented by - curves [30, 31, 32, 33]. For instance, Markakis et al. [34] provide correlations using numerical relativity simulations of the inspiral and postmerger GW signals from BNS mergers that can be used to constrain the EOS.
Agathos et al. [30], among the first to employ a Bayesian pipeline for EOS model selection, computed pairwise Bayes factors for each EOS and a collection of simulated BNS events. However, they approximated the tidal deformability of NSs by a Taylor expansion around a fiducial mass.
@@
In this study, we adopt the model selection framework introduced in Ref. [31] and extend its application to a population of simulated binary neutron star (BNS) signals expected from future gravitational-wave (GW) detectors with enhanced sensitivities. In addition, we choose to compute and compare the Bayesian evidence for different equations of state (EOSs) not only in the full four-dimensional parameter space defined by the component masses and their tidal deformabilities ---, but also within several lower-dimensional subspaces, namely , , and . By systematically evaluating the performance of EOS inference across these distinct subspaces, we are able to identify the most effective strategy for recovering the true EOS using the fourth model selection approach. We stress that this targeted choice of subspaces significantly improves the computational efficiency of the model selection procedure while preserving the essential information required to distinguish between different EOSs.
Given data which a model should explain, Bayesian evidence penalizes complex models and models with more parameters, technically known as Occam’s penalty. There are two known ways of enforcing Occam’s penalty. In the first case, one favors a model that conforms to a subjective prior before performing Bayesian inference. The other approach is naturally included in Bayesian inference, where the computation of evidence in a higher-dimensional space, corresponding to models with greater number of parameters, leads to a lower evidence [40, 41]. We incorporate the subjective prior by using EOS-dependent maximum mass limits through a reweighting of the posterior distribution that is initially EOS-agnostic and is obtained under the assumption of an arbitrarily large prior maximum mass [42]. Furthermore, for the evidence calculation we adopt an astrophysically motivated, reduced-dimensional parameter space, which helps ensure that any loss in evidence remains minimal when compared to carrying out the computation in the full parameter space.
@@
We defer the study of the impact of eccentricity to future analyses. Roy and Saini 56 has talked about the impact of eccentricity on the tidal deformability and concluded that it could have a large impact for BNS events observed by the XG detectors when the signal is analyzed from 10 Hz. We assume NSs to have finite, but small, spins which implies spin-induced quadrupole moments are sub-dominant effects in the waveform models. We have not looked at the impact of waveform uncertainties on our model selection results. We refer to Refs. [57, 43, 58, 20] for such discussions. Reali et al. 59 presents biases due to astrophysical confusion noise and found significant biases even for high signal-to-noise ratio (SNR) BNS mergers observed in next-generation detectors (however, also see [60] which finds the effect of foreground noise on parameter estimation to be minimal). It has also been argued that NS oscillations with timescales of the order of binary period could lead to biased estimate of NS tidal deformability and hence the EOS [61, 62, 63]. They mention the relative bias to be maximum of 50% for intermediate mass NS which is within the statistical uncertainty for current and near future configurations of GW detectors.
@@ II.1 Bayesian Evidence Calculation using EOS-agnostic PE results
where is the dimensionality of the prior parameter space which in our case is 4 for the case of BNS mergers. The total number of parameter for GW inference of BNS mergers has 17 parameters but, in our calculations, we start with the posterior PDF which is marginalized along all parameters except two masses () and two tidal deformabilities ().
@@
In our analysis, we derive the prior and posterior density estimates utilizing prior and posterior samples correspondingly. For each space, the samples are first transformed to another space where the posterior density is then computed. This is akin to applying transformations with Jacobians and additional marginalization, especially when integration is carried out in a lower-dimensional space.
The factor addresses the variance in effective prior volume between our analysis and the EOS-agnostic parameter estimation (PE) approaches. This difference emerges as our method enforces a narrow prior on the chirp mass to adhere to the specifications of the relative binning method in bilby. Hence, the evidence integration range in our study is considerably smaller than that in the EOS-agnostic PE, except when the event is near the maximum mass stipulated by a specific EOS. As a result, deviates noticeably from unity (typically ) for only a limited number of events, generally less than 10%, where the maximum allowable mass by the EOS is close to the observed posterior support. For the majority of events, this difference remains minimal. We have confirmed that this slight deviation of from 1.0 holds no substantial influence on our findings, since it impacts only a small portion of the sample and does not affect the overall trends of the Bayes factor.
@@ II.2 Combining Evidence from multiple events
where is the EOS-specific component mass prior correction factor taken into account differently for each EOS. We take this to be a uniform distribution between and the maximum mass allowed by that specific EOS. The factor is there to undo the prior used for the calculation of evidence for each event with the following value: for the calculation of 3D and 4D evidence and for the calculation of 2D evidence in the subspace. We keep our integration range slightly outside the range used in Bayesian PE run (except for the samples beyond the maximum mass of EOS), to allow for the kernel density estimation (KDE) to keep a conservative view of EOS-specific prior correction factor. We compared the KDE with the underlying sample distribution and found an excellent agreement (see Fig.13), effectively ruling out the risk of errors arising from the choice of density estimation method.
@@ II.3.1 Choice of EOS
Formally, to compare two models, it is essential to measure the difference between them, which has traditionally been referred to by the names of EOSs, highlighting the fundamental assumptions in the construction of EOSs. Although such concepts of distinguishability have significance, they are not useful in quantifying their differences. We formally quantify the disparity between different EOSs as the distance between them (refer to Eqn. (14) of Kashyap et al. [55] and B), with the caveat that this metric does not entirely capture their differences. In addition, other distance measures and correlations with populations will be crucial considerations, which we aim to explore in our upcoming research.
@@ III.2 Distribution of evidence
Degeneracies are anticipated to represent a fundamental limitation in differentiating distinct EOS models, as two different - relationships may yield identical observed posterior data. We find that the ambiguity in model selection due to the broad posterior distributions (due to the large noise in O5 and A) is greater than the apparent problem of degeneracy described above. Hence, we expect that will not be able to distinguish models in future GW detectors (such as ECC). However, we find that such differences are negligible and are important only when we combine the evidence from several events. As seen in Fig. 7, the degeneracy in is shown where two EOSs compete with the injected EOS, SLy in this case, up to a large number of events. However, the subspace evidence does slightly better.
@@ III.3 Cumulative Evidence from multiple BNS events
Additional clarity on detector sensitivity can be deduced from cumulative evidence plots. For example, the Adetector configuration is more capable of distinguishing a cluster of EOS that are similar to each other but significantly different from the injected EOS compared to O5 (Fig. 5). In the case of optimistic ECC observatories, the evidence for some of the (noninjected) EOS models are numerically zero, as shown in Fig. 8.
@@ III.4 Criterion for distinguishability and Method Efficacy
Moreover, we find that ignoring the pairwise correlation between and significantly improves the selection of the correct model. This is because such correlations do not contain information about the EOS and ignoring this i.e. avoid using the joint posterior, gives us the evidence in a smaller parameter space where Bayesian evidence calculations do not suffer from Occam’s penalty. This approach offers a new lens through which we can view and understand the interconnections between different methods proposed in the literature on model selection using the parameter . We emphasize that some correlations between BNS parameters are of astrophysical origin and are not determined by the EOS of NS matter, such as between mass and spin and between component masses. In addition, we point out that our current work is done for low-spin systems and hence mass-spin correlation measurement will be limited. We do not ignore the correlation between mass and tidal deformability, which is precisely what we aim to use and measure in our work. We aim to address correlations of astrophysical origin in our future research endeavors.
@@ IV Conclusions and Summary
For detectors with O5 sensitivity, evidence calculations across parameter space reveal no meaningful differences between the models in question. However, for the Aand ECC networks, performing evidence calculations in the subspace - is more effective in identifying the correct model than in other spaces. This is because excluding the mass ratio from evidence calculation does not eliminate any details concerning the EOS of nuclear matter. Additionally, this choice reduces the dimensional complexity of the evidence calculation.
@@ Appendix A Evidence distribution for APR4 and SLy
Appendix B Definition of
The -distance between the mass-radius or mass- curves can be used as a measure of the distance between a pair of EoS models. We propose that two models are more easily distinguishable greater is the -distance between the corresponding curves defined as:
where is normalization constants to render the distances dimensionless chosen to be:
and are the mass- curves corresponding to model and respectively, is the smallest NS mass in the observed population and is the smaller of the maximum mass allowed by models and
Appendix C KDE Reconstruction check